Simplify the following expression. Your answer should not contain logarithms.

25^log5(x)
(that's log base 5 then x)

Answer by venugopalramana(3286)

LET Y=25^(LOG X TO BASE 5)
USE FORMULA LOG X TO BASE 5 = LOG X/LOG 5 TO COMMON BASE
Y=25^(LOG X/LOG 5)
TAKING LOGS BOTH SIDES
LOG Y ={(LOG X)/(LOG 5)}*LOG 25
=LOG X*LOG 5^2/LOG 5=LOG X*(2*LOG 5)/LOG 5(SINCE LOG X^N=N*LOG X)
=2LOG X=LOG X^2
HENCE LOG Y = LOG X^2
TAKING ANTILOGS
Y=X^2
HENCE
Y=25^(LOG X TO BASE 5)=X^2